Find the constant of variation k
m varies directly with n; m=144 when n=8.

Answer by stanbon(48558)

Find the constant of variation k
m varies directly with n; m=144 when n=8.
varies directly means m=kn
To find "k" substitute for m and n to get:
144 = k(8)
k=144/8=18
Cheers,
stan H.